OpenStudy (anonymous):

Simplify: 2squared root32x^2y^3-xysquared rot 98y

OpenStudy (anonymous):

Does it not simplify?

OpenStudy (allank):

Kindly use the equation writing tool..

OpenStudy (anonymous):

\[2\sqrt{32x ^{2}y ^{3}}-xy \sqrt{98y}\]

OpenStudy (campbell_st):

well the 1st radical can be written as \[2\sqrt{16\times x^2\times y^2\times 2\times y}=2\sqrt{16x^2y^2}\times \sqrt{2y}\] which can be written as \[2 \times 4xy \times \sqrt{2y}\] the other radical is \[xy \times \sqrt{49\times2y}\] or \[7xy \sqrt{2y}\] so now after some simplification \[8xy \sqrt{2y} - 7xy \sqrt{2y}\] so you are subtracting like terms... you just need to finish it